OpenStudy (anonymous):

PLZZZZZZZZZZZZZZZZZZZZZZZZZZZZZZZZ HELP ME PLZ ILL GIVE A MEDAL PLZZZZZZZZZZZZZZ

OpenStudy (anonymous):

The temperature of a substance in a laboratory beaker dropped thirteen and one-half degrees Fahrenheit from its starting temperature, ending at sixty-eight and three-fourths degrees Fahrenheit. Let s equal the starting temperature. Which mathematical equation could be used to determine the starting temperature of the substance? s plus thirteen and one-half equals sixty-eight and three-fourths s plus sixty-eight and three-fourths equals thirteen and one-half s minus thirteen and one-half equals sixty-eight and three-fourths s times sixty-eight and three-fourths equals thirteen and one-half

OpenStudy (raden):

what's your question, plzzzzzzzzzz. lol

OpenStudy (anonymous):

Letting s be the starting temperature, "The temperature ... dropped thirteen and one-half degrees Fahrenheit from its starting temperature" So, \[s-13\frac{1}{2}\] "...ending at sixty-eight and three-fourths degrees Fahrenheit." Therefore, the equation you get is\[s-13\frac{1}{2}=68\frac{3}{4}\] Which one of the choices match that equation?